# SmallToLarge (S2L): Scalable Data Selection for Fine-tuning Large Language Models by Summarizing Training Trajectories of Small Models

### v1

@@ 5.4 Ablation Studies
[Uncaptioned image] Figure 7: S2L is robust to the length of training trajectories.
[Uncaptioned image] Figure 8: S2L prefers dense trajectories over sparse ones.
@@ 6 Conclusion and Limitations
Our study does come with its limitations. S2L has been only tested within two domains, mathematics and medicine, and on models up to 3 billion parameters, constrained by our computational resources. Additionally, our experiments employed a fixed training schedule across all methods without further optimization or hyperparameter tuning for each method, including S2L. This unified approach, while it ensures a fair comparison, may not fully capture the potential performance improvement that could be achieved with more tailored training strategies. We encourage further research to extend the application of S2L across a broader spectrum of domains and investigate the impact of hyperparameter tuning on its effectiveness.
Impact Statements
Appendix A Experiment Details
A.1 Models
@@
The Pythia models (Biderman et al., 2023b) are a suite of large language models (LLMs) developed by EleutherAI licensed under the Apache License 2.0. These models range in size from 70 million to 12 billion parameters and are designed to enable controlled scientific research on transparently trained LLMs across various scales.
@@
The Phi models (Li et al., 2023b) developed by Microsoft are under the MIT License. Phi-1.5, a transformer-based model with 1.3 billion parameters, and its successor, Phi-2, with 2.7 billion parameters, have been trained on a diverse set of data sources, including synthetic texts and curated websites. The Phi models underscore the potential of small yet powerful language models in understanding and generating human language, empowering a range of NLP tasks. Phi-2, in particular, has raised the bar for reasoning and language understanding among foundation models, matching or even exceeding the performance of models 25 times its size on complex benchmarks.
A.2 Datasets
@@
The MathInstruct dataset (Yue et al., 2023) is compiled from 13 diverse math rationale datasets, using both chain-of-thought (CoT) and program-of-thought (PoT) rationales. It ensures comprehensive coverage across various mathematical fields in the 262K training examples, making it a popular resource for fine-tuning large language models (LLMs) for general math problem-solving. MathInstruct is licensed under the MIT license.
@@
The MIMIC-III (Medical Information Mart for Intensive Care III) dataset (Johnson et al., 2016) is a comprehensive collection of de-identified health data associated with over 40,000 patients who stayed in critical care units of the Beth Israel Deaconess Medical Center in Boston, Massachusetts. This large dataset includes information such as demographics, vital signs, laboratory tests, medications, and more, making it an invaluable resource for a wide range of research in healthcare, including clinical decision support systems, medical procedure efficacy studies, and patient care optimization strategies.
@@
A.3 Implementation Details
@@
The training trajectories for both MathInstruct and MIMIC-III are gathered from training a Pythia-70M model, the smallest model in the Pythia model suite, recorded every $500$ training iterations. We utilize the Faiss library (Douze et al., 2024) to perform efficient K-means clustering of loss trajectories with Euclidean distance with $K=100$ and $20$ iterations. The hyperparameter $K$ is tuned in the range of $\{50,100,200\}$ based on the average accuracy of the model trained on $30K$ selected data. We found $K=100$ worked the best for both datasets we studied in this paper. Ablations studies on the length and the best time in the training to record the trajectories can be found in Section 5.4.
@@
For one-shot selection methods (excluding S2L), we use representations from either step 1000 or the end of fine-tuning Pythia-410M on MathInstruct and reported the better result in Table 1. In Table 4, we include the complete comparison between using early-fine-tuning vs. end-of-fine-tuning model checkpoints as the inference model. For Facility Locations, we further compared using the first hidden states as the feature representation as suggested in (Bhatt et al., 2024) to using the last hidden states (Wu et al., 2023b) for the tasks we studied.
Table 4: Complete results used for selecting the best reference model for each one-shot data selection baseline. The choice of early-fine-tuning (step 1000) and end-of-fine-tuning checkpoint follows (Marion et al., 2023). The best results selected for Table 1 are highlighted in cyan.
@@ A.4.1 MathInstruct
GSM8K
1319
Early Algebra
Multi-step reasoning using basic arithmetic operations
MATH
5000
Early Algebra, Intermediate Algebra, Algebra, Probability, NumTheory, Calculus, Geometry
Problems from mathematics competitions, including the AMC 10, AMC 12, AIME
NumGLUE
1042
Early Algebra
Commonsense, Domain-specific, Arithmetic Reasoning, Quantitative Comparison, Fill-in-the-blanks Format, Reading Comprehension, Numerical Reasoning, Quantitative NLI, Arithmetic Word Problems
SVAMP
1000
Early Algebra
Arithmetic Word Problems
Mathematics
1000
Early Algebra, Intermediate Algebra, NumTheory, Calculus
Arithmetic Reasoning
SimulEq
514
Linear Algebra
Single and multiple equation word problems
@@
We utilize the pipeline provided by (Yue et al., 2023), designed to first determine whether the model can be prompted to generate a code snippet. This code snippet, if successfully generated, should be executable and produce the correct answer when run. This code-based evaluation is also used for Phi models (Li et al., 2023b). In cases where the model does not directly produce a viable code solution, we employ a “think step-by-step” prompting strategy (Wei et al., 2022b). This method prompts the model to break down its reasoning process, a technique that has been widely proven effective in fully exploiting the model’s problem-solving capacity.
A.4.2 MIMIC-III
Following (Delbrouck et al., 2023; Demner-fushman et al., 2023), we include the six most common modality/anatomy pairs: CT head, CT abdomen, CT chest, MRI head, CT spine, and CT neck, and five less common pairs in the text data: MRI spine, CT sinus, MRI abdomen, MRI pelvis, and MRI neck in the evaluation. There are in total 13.7K test examples after data preprocessing and train-test splitting.
Appendix B Examples in Different Clusters
We provide the full version of Figure 5, which compares data points in the same and different clusters based on training trajectories, in Figure 9, Figure 10 and Figure 11. We can observe that examples with similar training trajectories have the same question format. Therefore, balanced sampling from all clusters can ensure different types of examples can be covered in the selected subset of training data.
Figure 9: A full version of Figure 4(a): open-formed algebra. Questions are in black and answers are in cyan.
Figure 10: A full version of Figure 4(b): reading comprehension + coding. Questions are in black and answers are in cyan; instructions are highlighted in orange.
Figure 11: A full version of Figure 4(c): multiple-choice + multi-step reasoning. Questions are in black and answers are in cyan; instructions are highlighted in orange.
Appendix C Topic Distribution of Data Selected by S2L

### v2

@@ 5.4 Ablation Studies
Figure 8: S2L is robust to the length of training trajectories.
Figure 9: S2L prefers dense trajectories over sparse ones.
@@
S2L does not require the full training data to train the proxy and can scale efficiently to larger datasets. To further demonstrate the scalability of the proposed S2L method, we conducted experiments by training the proxy on a smaller sample of the data (100K examples) for the same number of epochs (3 epochs) and saving the loss for all examples. The results, shown in Figure 11, confirm that S2L remains effective when the proxy model is trained on a smaller subset of training data and therefore is scalable to larger datasets without a proportional increase in computational costs.
S2L is robust across different clustering parameter values for K. We conducted detailed experiments varying the clustering parameter K, as shown in Figure 11. The results demonstrate that S2L maintains high performance across different values of K, highlighting the robustness of our method to different clustering parameter choices. We chose K=100 for our experiments as it provided the best average accuracy across the evaluation datasets for the math reasoning task.
S2L remains effective and efficient compared to using full data when trained for the same number of epochs. Figure 12 illustrates the relative accuracy to full data across different epochs, comparing S2L-selected data and full data with the same number of epochs. Both in-domain and overall average accuracy are shown. S2L demonstrates superior performance with fewer data and fewer training iterations.
S2L supports a range of small models as effective proxies. To understand whether different small models could serve as effective proxies, we used GPT-2 (124M) and Pythia-160M as proxy models for data selection to train Pythia-410M. The results, illustrated in Figure 13, show that both proxy models perform comparably in guiding the data selection, demonstrating the versatility and effectiveness of using different small models for S2L.
Figure 10: Per-dataset and average accuracy comparing proxy training on 100K examples and full data. S2L remains effective.
Figure 11: Per-dataset and average accuracy across different values of the clustering parameter $K$. S2L is relatively robust to the choice of $K$.
(a) In-domain Average Accuracy
(b) Overall Average Accuracy
Figure 12: Relative accuracy to full data across different epochs, comparing S2L-selected data and full data. S2L achieves superior performance with fewer data and fewer training iterations.
Figure 13: Per-dataset and average accuracy comparison between using different proxy models (Pythia-160M and GPT-2 (124M)) for data selection. Using both proxy models show comparable performance, demonstrating the effectiveness of different small models as reference models for S2L.
@@ 6 Conclusion and Limitations
Our study does come with its limitations. S2L has been only tested within two domains, mathematics and medicine, and on models up to 7 billion parameters, constrained by our computational resources. Additionally, our experiments employed a fixed training schedule across all methods without further optimization or hyperparameter tuning for each method, including S2L. This unified approach, while it ensures a fair comparison, may not fully capture the potential performance improvement that could be achieved with more tailored training strategies. We encourage further research to extend the application of S2L across a broader spectrum of domains and investigate the impact of hyperparameter tuning on its effectiveness.
Acknowledgments
This research was partially supported by the National Science Foundation CAREER Award 2146492, National Science Foundation 2421782 and Simons Foundation, Cisco Systems, Optum AI, and a UCLA Hellman Fellowship.
Appendix A Proofs
A.1 Proof of Theorem 4.1
Proof.
From the assumption that the loss trajectories of examples on the proxy and target models are close:
$\|\mathbf{L}_{i}^{\text{proxy}}-\mathbf{L}_{i}^{\text{target}}\|\leq\delta,\quad\forall i.$ (9)
Since $i$ and $j$ are in the same cluster $C_{k}$ based on the proxy model, we have:
$\|\mathbf{L}_{i}^{\text{proxy}}-\mathbf{L}_{j}^{\text{proxy}}\|\leq\epsilon.$ (10)
Using the triangle inequality:
$\|\mathbf{L}_{i}^{\text{target}}-\mathbf{L}_{j}^{\text{target}}\|\leq\|\mathbf{L}_{i}^{\text{target}}-\mathbf{L}_{i}^{\text{proxy}}\|+\|\mathbf{L}_{i}^{\text{proxy}}-\mathbf{L}_{j}^{\text{proxy}}\|+\|\mathbf{L}_{j}^{\text{proxy}}-\mathbf{L}_{j}^{\text{target}}\|\leq 2\delta+\epsilon=\epsilon^{\prime}.$ (11)
Therefore, at any iteration $t$:
$|\mathcal{L}_{i}^{\text{target}}(\bm{\theta}^{(t)})-\mathcal{L}_{j}^{\text{target}}(\bm{\theta}^{(t)})|\leq\epsilon^{\prime},\quad\forall t.$ (12)
Assuming that the loss functions can be approximated by:
$\mathcal{L}_{i}^{\text{target}}(\bm{\theta})=\frac{1}{2}d\bm{\theta}^{\top}\bm{H}_{i}d\bm{\theta}+\bm{g}_{i}^{\top}d\bm{\theta}+c_{i},$ (13)
where $c_{i}$ is the loss of example $i$ at the beginning of fine-tuning, and $d\bm{\theta}$ is the distance between the parameters of the pretrained model and those during fine-tuning. Similarly for $\mathcal{L}_{j}^{\text{target}}(\bm{\theta})$. The loss difference between $i$ and $j$ is:
$\mathcal{L}_{i}^{\text{target}}(\bm{\theta})-\mathcal{L}_{j}^{\text{target}}(d\bm{\theta})=\frac{1}{2}d\bm{\theta}^{\top}(\bm{H}_{i}-\bm{H}_{j})d\bm{\theta}+(\bm{g}_{i}-\bm{g}_{j})^{\top}d\bm{\theta}+(c_{i}-c_{j}).$ (14)
Given that $|\mathcal{L}_{i}^{\text{target}}(\bm{\theta})-\mathcal{L}_{j}^{\text{target}}(\bm{\theta})|\leq\epsilon^{\prime}$, we can write:
$\left|\frac{1}{2}d\bm{\theta}^{\top}(\bm{H}_{i}-\bm{H}_{j})d\bm{\theta}+(\bm{g}_{i}-\bm{g}_{j})^{\top}d\bm{\theta}+(c_{i}-c_{j})\right|\leq\epsilon^{\prime}.$ (15)
Let us choose two different values, $\bm{\theta}^{(1)}$ and $\bm{\theta}^{(2)}$, to generate two inequalities. For $d\bm{\theta}^{(1)}$, we have:
$\left|\frac{1}{2}(d\bm{\theta}^{(1)})^{\top}(\bm{H}_{i}-\bm{H}_{j})d\bm{\theta}^{(1)}+(\bm{g}_{i}-\bm{g}_{j})^{\top}d\bm{\theta}^{(1)}+(c_{i}-c_{j})\right|\leq\epsilon^{\prime},$ (16)
and for $d\bm{\theta}^{(2)}$, we have:
$\left|\frac{1}{2}(d\bm{\theta}^{(2)})^{\top}(\bm{H}_{i}-\bm{H}_{j})d\bm{\theta}^{(2)}+(\bm{g}_{i}-\bm{g}_{j})^{\top}d\bm{\theta}^{(2)}+(c_{i}-c_{j})\right|\leq\epsilon^{\prime}.$ (17)
Subtracting these two inequalities, we get:
$\left|\frac{1}{2}\left((d\bm{\theta}^{(1)})^{\top}(\bm{H}_{i}-\bm{H}_{j})\bm{\theta}^{(1)}-(d\bm{\theta}^{(2)})^{\top}(\bm{H}_{i}-\bm{H}_{j})d\bm{\theta}^{(2)}\right)+(\bm{g}_{i}-\bm{g}_{j})^{\top}(d\bm{\theta}^{(1)}-d\bm{\theta}^{(2)})\right|\leq 2\epsilon^{\prime}.$ (18)
$\begin{split}\left|(d\bm{\theta}^{(1)})^{\top}(\bm{H}_{i}-\bm{H}_{j})d\bm{\theta}^{(1)}-(d\bm{\theta}^{(2)})^{\top}(\bm{H}_{i}-\bm{H}_{j})d\bm{\theta}^{(2)}\right|&\leq\|\bm{H}_{i}-\bm{H}_{j}\|\left(\|d\bm{\theta}^{(1)}\|^{2}+\|d\bm{\theta}^{(2)}\|^{2}\right)\\ &\leq(\|\bm{H}_{i}\|+\|\bm{H}_{j}\|)\left(\|d\bm{\theta}^{(1)}\|^{2}+\|d\bm{\theta}^{(2)}\|^{2}\right)\\ &\leq 4CD^{2}\end{split}$ (19)
This gives us:
$\left|(\bm{g}_{i}-\bm{g}_{j})^{\top}(d\bm{\theta}^{(1)}-d\bm{\theta}^{(2)})\right|\leq 2\epsilon^{\prime}+2CD^{2}.$ (20)
Assuming $\|d\bm{\theta}^{(1)}-d\bm{\theta}^{(2)}\|\geq d$, we get:
$\|\bm{g}_{i}-\bm{g}_{j}\|\leq\frac{2\epsilon^{\prime}+2CD^{2}}{d}=\Delta.$ (21)
A.2 Proof of Corollary 4.2
Without loss of generality, assume we select $k$ example from each cluster and we have $k\leq\min_{j\in[K]}|C_{j}|$. Then the error of the subset in capturing the full gradient will be
$\xi\leq\sum_{j}(|C_{j}|-k)(\bar{\bm{g}}_{j}+\Delta),$ (22)
where $\bar{\bm{g}}_{j}$ is the norm of the average gradient of the selected examples from $C_{j}$. In practice, we can weight elements of the subset by $r_{\min}/k$, which has a similar effect to scaling the step size when training on the subset. Let ${\bm{g}}_{\max}=\max_{j}\|{\bm{g}}_{j}\|$ be the maximum gradient norm during training, $r_{\max}=\max_{j}|C_{j}|,r_{\min}=\min_{j}|C_{j}|$. Then, we get
$\xi^{\prime}$ $\leq\sum_{j}(r_{\min}-k)\Delta+(|C_{j}|-r_{\min})(\bar{\bm{g}_{j}}+\Delta)$ (23)
$\leq K[r_{\min}\Delta+(r_{\max}-r_{\min})\bm{g}_{\max}]$ (24)
The first term in RHS of Eq (23) is the error of the subset selected from $C_{j}$ to capture its full gradient and the second term is due to selecting the same number of examples, $k$, from the larger clusters.
Using the above error and following the proof of Theorem 1 in [35], for a constant step size $\alpha\leq 1/c$ we get:
$\|\bm{\theta}^{t+1}-\bm{\theta}^{*}\|^{2}\leq(1-\alpha c)^{t+1}\|\bm{\theta}^{t}-\bm{\theta}^{*}\|^{2}+2\xi^{\prime}R/c^{2}+\alpha B^{2}(r_{\min}/k)^{2}\bm{g}_{\max}^{2},$ (25)
where $c\leq\|\bm{H}\|$, and $B=k\cdot K$ is the total size of the subset, $R=\min\{d_{0},B\bm{g}_{\max}+\xi^{\prime}/c\}$ and $d_{0}=\|\bm{\theta}^{0}-\bm{\theta}^{*}\|$ is the initial distance to the optimal solution $\bm{\theta}^{*}$.
If $k\geq|C_{j}|$ for any cluster $C_{j}$, one can simply add $(r_{\min}/k-1)\cdot\hat{\bm{g}}_{j}$ to $\xi^{\prime}$ for the corresponding clusters, where $\hat{\bm{g}_{j}}$ is the norm of the total gradient of cluster $C_{j}$ and we replace $r_{\min}$ in Eq (23) with the size of smallest cluster that has larger than $k$ examples.
Appendix B Experiment Details
B.1 Models
@@
The Pythia models [6] are a suite of large language models (LLMs) developed by EleutherAI licensed under the Apache License 2.0. These models range in size from 70 million to 12 billion parameters and are designed to enable controlled scientific research on transparently trained LLMs across various scales.
@@
The Phi models [28] developed by Microsoft are under the MIT License. Phi-1.5, a transformer-based model with 1.3 billion parameters, and its successor, Phi-2, with 2.7 billion parameters, have been trained on a diverse set of data sources, including synthetic texts and curated websites. The Phi models underscore the potential of small yet powerful language models in understanding and generating human language, empowering a range of NLP tasks. Phi-2, in particular, has raised the bar for reasoning and language understanding among foundation models, matching or even exceeding the performance of models 25 times its size on complex benchmarks.
LLaMA 2.
The LLaMA 2 models [50], released by Meta AI and licensed under the LLaMA 2 Community License Agreement, are designed for improved natural language understanding and generation. LLaMA 2-7B, the smallest in this series with 7 billion parameters, has demonstrated competitive performance across various NLP benchmarks despite its moderate size.
B.2 Datasets
@@
The MathInstruct dataset [64] is compiled from 13 diverse math rationale datasets, using both chain-of-thought (CoT) and program-of-thought (PoT) rationales. It ensures comprehensive coverage across various mathematical fields in the 262K training examples, making it a popular resource for fine-tuning large language models (LLMs) for general math problem-solving. MathInstruct is licensed under the MIT license.
@@
The MIMIC-III (Medical Information Mart for Intensive Care III) dataset [21] is a comprehensive collection of de-identified health data associated with over 40,000 patients who stayed in critical care units of the Beth Israel Deaconess Medical Center in Boston, Massachusetts. This large dataset includes information such as demographics, vital signs, laboratory tests, medications, and more, making it an invaluable resource for a wide range of research in healthcare, including clinical decision support systems, medical procedure efficacy studies, and patient care optimization strategies.
@@
Table 3: A synthetic radiology report (MRI of the brain), generated by the GPT-4 model [2] to demonstrate the typical data format and content used in the clinical text summarization task. It is not suitable for clinical or diagnostic use.
Findings The brain parenchyma demonstrates normal morphology with no evidence of mass effect or midline shift. No acute infarcts are seen on diffusion-weighted images. There are no signs of intracranial hemorrhage. Mild generalized cerebral atrophy is noted. The ventricles and sulci appear within normal limits for the patient’s age. The pituitary gland and sella turcica are unremarkable. There are no abnormal signal intensities within the brain parenchyma. The orbits, paranasal sinuses, and mastoid air cells are clear.
Impression Normal MRI of the brain. Mild cerebral atrophy, likely age-related. No acute intracranial pathology.
B.3 Implementation Details
@@
The training trajectories for both MathInstruct and MIMIC-III are gathered from training a Pythia-70M model, the smallest model in the Pythia model suite, recorded every $500$ training iterations. We utilize the Faiss library [17] to perform efficient K-means clustering of loss trajectories with Euclidean distance with $K=100$ and $20$ iterations. The hyperparameter $K$ is tuned in the range of $\{50,100,200\}$ based on the average accuracy of the model trained on $30K$ selected data. We found $K=100$ worked the best for both datasets we studied in this paper. Ablations studies on the length and the best time in the training to record the trajectories can be found in Section 5.4.
@@
For one-shot selection methods (excluding S2L), we use representations from either step 1000 or the end of fine-tuning Pythia-410M on MathInstruct and reported the better result in Figure 4 and Table 1. In Table 4, we include the complete comparison between using early-fine-tuning vs. end-of-fine-tuning model checkpoints as the inference model. For Facility Locations, we further compared using the first hidden states as the feature representation as suggested in [4] to using the last hidden states [56] for the tasks we studied.The ranges for confidence, perplexity, and learnability are chosen according to the best-performing intervals reported in prior research (Section 5.1).
Due to memory and computational constraints, for Facility Locations, we calculate pairwise similarity and perform greedy selection on a per-data-source basis. We found this per-source selection approach also yields benefits for S2L as different data sources within MathInstruct exhibit distinct common patterns in their training trajectories. Therefore, we implement S2L also on a per-source basis for MathInstruct, and recommend applying S2L per source when dealing with datasets composed of multiple data sources.
Hyperparameters
Following the setup used in [64], we adopt a training regimen with a learning rate of 2e-5, a batch size of 128, a maximum length of 512, and a cosine scheduler with a 3% warm-up period.
Experiments Compute Resources
We fine-tune all the models with the Huggingface transformers library [55] with Fully Sharded Data Parallel (FSDP) [66] on 4 48G NVIDIA RTX A6000.
Table 4: Complete results used for selecting the best reference model for each one-shot data selection baseline. The choice of early-fine-tuning (step 1000) and end-of-fine-tuning checkpoint follows [34]. The best results selected for Figure 4 are highlighted in cyan.
@@ B.4.1 MathInstruct
GSM8K 1319 Early Algebra Multi-step reasoning using basic arithmetic operations
MATH 5000 Early Algebra, Intermediate Algebra, Algebra, Probability, NumTheory, Calculus, Geometry Problems from mathematics competitions, including the AMC 10, AMC 12, AIME
NumGLUE 1042 Early Algebra Commonsense, Domain-specific, Arithmetic Reasoning, Quantitative Comparison, Fill-in-the-blanks Format, Reading Comprehension, Numerical Reasoning, Quantitative NLI, Arithmetic Word Problems
SVAMP 1000 Early Algebra Arithmetic Word Problems
Mathematics 1000 Early Algebra, Intermediate Algebra, NumTheory, Calculus Arithmetic Reasoning
SimulEq 514 Linear Algebra Single and multiple equation word problems
@@
We utilize the pipeline provided by [64] [^2] https://github.com/TIGER-AI-Lab/MAmmoTH?tab=readme-ov-file#large-scale-evaluation, designed to first determine whether the model can be prompted to generate a code snippet. This code snippet, if successfully generated, should be executable and produce the correct answer when run. This code-based evaluation is also used for Phi models [28]. In cases where the model does not directly produce a viable code solution, we employ a “think step-by-step" prompting strategy [54]. This method prompts the model to break down its reasoning process, a technique that has been widely proven effective in fully exploiting the model’s problem-solving capacity.
B.4.2 MIMIC-III
Following [14, 15], we include the six most common modality/anatomy pairs: CT head, CT abdomen, CT chest, MRI head, CT spine, and CT neck, and five less common pairs in the text data: MRI spine, CT sinus, MRI abdomen, MRI pelvis, and MRI neck in the evaluation. There are in total 13.7K test examples after data preprocessing and train-test splitting.
(a) GSM8K
(b) MATH
(c) NumGLUE
(d) SVAMP
(e) Mathematics
(f) SimulEq
(g) In-domain Avg
(h) Avg
Figure 14: Break-down accuracies ($\uparrow$) on in-domain and out-of-domain datasets using Pythia-410M. Data size refers to the total number of unique training examples used. All experiments in this table share the same total training steps and learning rate schedule (see Section 5.2).
Appendix C Examples in Different Clusters
We compare data points in the same and different clusters based on training trajectories, in Figure 15, Figure 16 and Figure 17. We can observe that examples with similar training trajectories have the same question format. Therefore, balanced sampling from all clusters can ensure different types of examples can be covered in the selected subset of training data.
Figure 15: Examples in the cluster shown in Figure 7(a): open-formed algebra. Questions are in black and answers are in cyan.
Figure 16: Examples in the cluster shown in Figure 7(b): reading comprehension + coding. Questions are in black and answers are in cyan; instructions are highlighted in orange.
Figure 17: Examples in the cluster shown in Figure 7(c): multiple-choice + multi-step reasoning. Questions are in black and answers are in cyan; instructions are highlighted in orange.
Appendix D Topic Distribution of Data Selected by S2L